A rectangular park is 95 meters long and 50 meters wide.
Give the length and width of another rectangular park that has the same perimeter but a smaller area.
What's the length and width???

Answer by richwmiller(17219)

2*95+2*50=290
2w+2l=290
l+w=145
l*w<4750
There are 98 integer solutions
l = 1, w = 144
l = 2, w = 143